the is the formula of tan3A...................................................???????????????

Grade:12

13 Answers

NAVEENKUMAR NAGIPOGU
8 years ago
dear mohan,
      the formula of tan3a is the ratio of  3TANA-TAN ^3A – 1-TAN^2A.THIS IS ONLY FROM THE KNOWN AND TRUTH.FOR MORE DETAILS AND PROOVES AGAIN ASK.
N JYOTHEESWAR
8 years ago
writing tan(3A) as tan(2A + A) and use the expression for tan(A + B) to get. tan(2A + A) multiply with tan (a+b).we get tan 3a.
Prabhakar ch
8 years ago
Dear Mohan
                    Your answer is 3tana-tan^3a-1-tan^2a
All the best for your bright future...........................
Gowri sankar
8 years ago
HI MOHAN
                The formula of TAN3 IS...............................................................................................................3tana-tan^3a-1-tan^2a.
raj
8 years ago
hallo mohan Your answer is 3tana-tan^3a-1-tan^2a.............................................................
T Dileep
8 years ago
hai mohan,formula of tan3x=(3tanx-tan^3x)/(1-3tan^2x)this is the formula of tan3x.tan3x formula in also represents another form.
sudarshan
8 years ago
the formula of tan3a is the ratio of 3TANA-TAN^3A-1-TAN^2A.This is the formula of tan3a.................
T.kumar
8 years ago
tan3a=3tan-tan^3a /1-3tan^2a. this is my answer .. all the best foyr your global furture...please approve..
Y MANJUNADH
8 years ago
mohan
              the formula of   tan3a  is  (3tana-tan3a)/1-3tan2a............................................it is used in trigonometry
 
manichandra
8 years ago
hi, i...indeed,to..need..your..help how..cn..i..prove..this, formula??? sin3a=3sina-4sin^3a cos3a=4cos^3-3cosa tan3a=3tan-tan^3a /1-3tan^2a. the..symbol.
BALAJI ANDALAMALA
8 years ago
tan3A=tan(2A+A)=\frac{tan2A+tanA}{1-tan2A.tanA}
=\frac{\frac{2tanA}{1-tan^2A}+tanA}{1-\frac{2tanA}{1-tan^2A}.tanA}
=\frac{3tanA-tan^3A}{1-3tan^2A}
siva prasad reddy
8 years ago
 
Dear Mohan
                    Your answer is 3tana-tan^3a-1-tan^2a
All the best for your bright future...........................
 
thank you
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

tan3x=(3tanx-tan^3x)/(1-3tan^2x)

Thanks and Regards